# SPECIAL RELATIVITY

### v9

@@ Appendix I. Radiation by Moving Charges
The only assumption made going from Eq. (138) to Eq. (144) is that edge terms in the integration by parts can be dropped. This assumption can be justified by means of physical arguments in the most general situation accounting for the fact that the integral in $dt^{\prime}$ has to be performed over the entire history of the particle and that at $t^{\prime}=-\infty$ and $t^{\prime}=+\infty$ the electron does not contribute to the field anymore. Let us give a concrete example for an ultra-relativistic electron. Imagine that bending magnets are placed at the beginning and at the end of a given setup, such that they deflect the electron trajectory of an angle much larger than the maximal observation angle of interest for radiation from a bending magnet. This means that the magnets would be longer than the formation length associated with the bends, i.e. $L_{\mathrm{fb}}\sim(c\rho^{2}/\omega)^{1/3}$, where $\rho$ is the bending radius. In this way, intuitively, the magnets act like switches: the first magnet switches the radiation on, the second switches it off. Then, what precedes the upstream bend and what follows the downstream bend does not contribute to the field detected at the screen position. With these caveat Eq. (144) is completely equivalent to Eq. (138).
The derivation of Eq. (144) is particularly instructive because shows that each term in Eq. (144) is due to a combination of velocity and acceleration terms in Eq. (138). In other words the terms in $1/R$ and in $1/R^{2}$ in Eq. (144) appear as a combination of the terms in $1/R$ (acceleration term) and $1/R^{2}$ (velocity term) in Eq. (138). As a result, one can say that there exist contributions to the radiation from the velocity part in Eq. (138). The presentation in Eq. (144) is more interesting with respect to that in Eq. (138) (although equivalent to it) because the magnitude of the $1/R^{2}$-term in Eq. (144) can directly be compared with the magnitude of the $1/R$-term inside the integral sign.
@@
If the position of the observer is far away enough from the charge, one can make the expansion Eq. (30). Using Eq. (148), we obtain Eq. (LABEL:revwied).
Appendix II. Undulator Radiation in Resonance Approximation. Far Zone
@@
Within the resonance approximation we further select frequencies such that
$\frac{|\Delta\omega|}{\omega_{r}}\ll 1~{},~{}~{}~{}~{}\mathrm{i.e.}~{}~{}|C|\ll k_{w}~{}.$ (161)
Note that this condition on frequencies automatically selects observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$. In fact, if one considers observation angles outside the range $\theta^{2}\ll 1/\gamma_{z}^{2}$, condition (160) is not fulfilled, and the integrand in Eq. (158) exhibits fast oscillations on the integration scale $L$. As a result, one obtains zero transverse field, $\vec{\widetilde{E}}=0$, with accuracy $1/N_{w}$. Under the constraint imposed by (161), independently of the value of $K$ and for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, we have
@@
This means that, independently of $K$, $|v|\ll 1$ and we may expand $J_{n}(v)$ in Eq. (158) according to $J_{n}(v)\simeq[2^{-n}/\Gamma(1+n)]~{}v^{n}$, $\Gamma(\cdot)$ being the Euler gamma function
@@
Here we are dealing specifically with the first harmonic. Therefore, these extra-simplifications apply. We neglect both the term in $\cos(k_{w}z^{\prime})$ in the phase of Eq. (152) and the term in $\vec{\theta}$ in Eq. (152). First, non-negligible terms in the expansion of $J_{n}(v)$ are those for small values of $n$, since $J_{n}(v)\sim v^{n}$, with $|v|\ll 1$. The value $n=0$ gives a non-negligible contribution $J_{0}(v)\sim 1$. Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq. (158) but those for $m=-1$ and $m=0$ average to zero due to resonance approximation. Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq. (158) averages to zero for $n=0$. Values $n=\pm 1$ already give negligible contributions. In fact, $J_{\pm 1}(v)\sim v$. Then, the term in $\vec{e}_{x}$ in Eq. (158) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$. The term in $\vec{\theta}$ would survive averaging when $n=1,~{}m=-1$ and when $n=-1,~{}m=0$. However, it scales as $\vec{\theta}v$. Now, using condition (161) we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$. Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$. All terms corresponding to larger values of $|n|$ are negligible.
@@
Appendix III. Approximation for the Electron Path
@@
Appendix IV. Self-Electromagnetic Fields of the Modulated Electron Beam

### v10

@@ 21 Appendix
The only assumption made going from Eq. (127) to Eq. (133) is that edge terms in the integration by parts can be dropped. This assumption can be justified by means of physical arguments in the most general situation accounting for the fact that the integral in $dt^{\prime}$ has to be performed over the entire history of the particle and that at $t^{\prime}=-\infty$ and $t^{\prime}=+\infty$, the electron does not contribute to the field anymore. Let us give a concrete example for an ultra-relativistic electron. Imagine that bending magnets are placed at the beginning and at the end of a given setup, such that they deflect the electron trajectory of an angle much larger than the maximal observation angle of interest for radiation from a bending magnet. This means that the magnets would be longer than the formation length associated with the bends, i.e. $L_{\mathrm{fb}}\sim(c\rho^{2}/\omega)^{1/3}$, where $\rho$ is the bending radius. In this way, intuitively, the magnets act like switches: the first magnet switches the radiation on, and the second switches it off. Then, what precedes the upstream bend and what follows the downstream bend does not contribute to the field detected at the screen position. With this caveat Eq. (133) is completely equivalent to Eq. (127).
The derivation of Eq. (133) is particularly instructive because shows that each term in Eq. (133) is due to a combination of velocity and acceleration terms in Eq. (127). In other words the terms in $1/R$ and in $1/R^{2}$ in Eq. (133) appear as a combination of the terms in $1/R$ (acceleration term) and $1/R^{2}$ (velocity term) in Eq. (127). As a result, one can say that there exist contributions to the radiation from the velocity part in Eq. (127). The presentation in Eq. (133) is more interesting with respect to that in Eq. (127) (although equivalent to it) because of the magnitude of the $1/R^{2}$-term in Eq. (133) can directly be compared with the magnitude of the $1/R$-term inside the integral sign.
@@
If the position of the observer is far away enough from the charge, one can make the expansion $\left|\vec{r}_{0}-\vec{r}(t)\right|=r_{0}-\vec{n}\cdot\vec{r}(t)$. Using Eq. (137), we obtain Eq. (LABEL:revwied).
21.2 A2. Undulator Radiation in Resonance Approximation. Far Zone
@@
Within the resonance approximation we further select frequencies such that $|\Delta\omega|/\omega_{r}\ll 1$ (i.e. $|C|\ll k_{w}$). Note that this condition on frequencies automatically selects observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$. In fact, if one considers observation angles outside the range $\theta^{2}\ll 1/\gamma_{z}^{2}$, condition (149) is not fulfilled, and the integrand in Eq. (147) exhibits fast oscillations on the integration scale $L$. As a result, one obtains zero transverse field, $\vec{\widetilde{E}}=0$, with accuracy $1/N_{w}$. Under the constraint imposed by $|\Delta\omega|/\omega_{r}\ll 1$ , independently of the value of $K$ and for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, we have
@@
Here we are dealing specifically with the first harmonic. Therefore, these extra-simplifications apply. We neglect both the term in $\cos(k_{w}z^{\prime})$ in the phase of Eq. (141) and the term in $\vec{\theta}$ in Eq. (141). First, non-negligible terms in the expansion of $J_{n}(v)$ are those for small values of $n$, since $J_{n}(v)\sim v^{n}$, with $|v|\ll 1$. The value $n=0$ gives a non-negligible contribution $J_{0}(v)\sim 1$. Then, since the integration in $dz^{\prime}$ is performed over a large number of undulator periods $N_{w}\gg 1$, all terms of the expansion in Eq. (147) but those for $m=-1$ and $m=0$ average to zero due to resonance approximation. Note that surviving contributions are proportional to $K/\gamma$, and can be traced back to the term in $\vec{e}_{x}$ only, while the term in $\vec{\theta}$ in Eq. (147) averages to zero for $n=0$. Values $n=\pm 1$ already give negligible contributions. In fact, $J_{\pm 1}(v)\sim v$. Then, the term in $\vec{e}_{x}$ in Eq. (147) is $v$ times the term with $n=0$ and is immediately negligible, regardless of the values of $m$. The term in $\vec{\theta}$ would survive averaging when $n=1,~{}m=-1$ and when $n=-1,~{}m=0$. However, it scales as $\vec{\theta}v$. Now, using condition $|\Delta\omega|/\omega_{r}\ll 1$ we see that, for observation angles of interest $\theta^{2}\ll 1/\gamma_{z}^{2}$, $|\vec{\theta}|~{}|v|\sim(\sqrt{2}~{}K~{}/\sqrt{2+K^{2}}~{})~{}\bar{\gamma}_{z}\theta^{2}\ll K/\gamma$. Therefore, the term in $\vec{\theta}$ is negligible with respect to the term in $\vec{e}_{x}$ for $n=0$, that scales as $K/\gamma$. All terms corresponding to larger values of $|n|$ are negligible.
@@
21.3 A3. Approximation for the Electron Path
@@
21.4 A4. Self-Electromagnetic Fields of the Modulated Electron Beam